OpenStudy (anonymous):

A rock dropped from a cliff covers one-third of its total distance to the ground in the last second of its fall. Air resistance is negligible. How high is the cliff?

OpenStudy (anonymous):

@experimentX @shubhamsrg @JFraser

OpenStudy (jfraser):

this is already in the physics group

OpenStudy (anonymous):

Can u Check My answer i got h = 148.48 m

OpenStudy (experimentx):

s = 1/2 gt^2 ----------- v = 0 + g(t-1) now u = v then, 1/3 s = v(1) + 1/2 g(1)^2 s = 1/2 gt^2 ---------- from these two solve or 's' and 't' rest should be okay

OpenStudy (aravindg):

nicely explained @experimentX

OpenStudy (experimentx):

thanks!!

OpenStudy (anonymous):

Ok....i Did by this way: h = 1/2 g t^2 Dn = u + a/2 (2n-1) h/3 = a/2(2t-1).....?

OpenStudy (shubhamsrg):

in the formula you used @Yahoo! 'n' is the nth second, in your case, n= t-1 .. you already know h= 1/2 gt^2 also h/3 =g/2 (2t -3) taking g=10, i get h=45..someone please confirm ?

OpenStudy (anonymous):

But why Do we take t-1

OpenStudy (shubhamsrg):

last second right ? after t seconds are complete , the object doesnt travel further..

OpenStudy (anonymous):

I am Silly....u r correct...i got it now....

OpenStudy (anonymous):

thxxxx

OpenStudy (shubhamsrg):

i may be wrong in logic though.. let someone else make any correction if its there..

OpenStudy (anonymous):

I think ur Logic is Correct...

OpenStudy (anonymous):

But Lets Confirm..as u told..@sauravshakya

OpenStudy (shubhamsrg):

maybe not tagged correctly..allow me @sauravshakya

OpenStudy (shubhamsrg):

you got 148 right? by another plausible method, i am also getting 148, so that should be right,,my logic is thus, flawed//hmm..

OpenStudy (shubhamsrg):

here is what i did,, we h = 1/2 g t1 ^2 ....(1) where t1 is total time taken . let t2 be time before it comes to last second of the journey, thus t1 - t2 =1 => t2 = t1 -1....(2) we know h- h/3 = 1/2 g t2 ^2 2h/3 = 1/2 g t2 ^2 ...(3) using 1,2,and 3 eqns, we can solve for h, t1 and t2 .. by taking g=10, i approximately get h=148.4 something.. which should be the ans.. unless ofcorse,,there is some flaw in this also! :|

OpenStudy (anonymous):

This one looks correct

OpenStudy (anonymous):

But This one Luks Similar to 1st One @shubhamsrg

OpenStudy (shubhamsrg):

1st one ?